in sulphur estimation 0.157 gm of organic compound gave 0.4813gm of BaSO4. Calculate the % of S
1
Expert's answer
2012-09-14T09:15:02-0400
m 0.4813
S-----BaSO4
32 & & & 233

m = 32*0.4813/233 =
0.0661

w=0.0661/0.157 * 100%=42.1 %